# Adaptive Exploration for Data-Efficient General Value Function Evaluations

### v1

@@ 8.2 Continuous State Environment with Non-Linear Function Approximation
We use a Continuous GridWorld environment that extends the tabular experiments to a continuous state space (McLeod et al., 2021), and maintain four discrete cardinal actions. This environment is purposefully selected for its suitability in demonstrating meaningful and distinct GVF goals, offering a clearer qualitative analysis compared to more complex environments. We examine Setting 4 and consider two GVFs. The first GVF has a cumulant at the top-left corner $c_{1}=\mathcal{N}(\mu=100,\sigma=10)$ and the second GVF has at the top-right corner $c_{2}=\mathcal{N}(\mu=50,\sigma=5)$, with zero cumulant assigned everywhere else. We used a constant discount $\gamma=0.99$. We used an Experience Replay Buffer with capacity $25K$ and batch size $64$ for all the experiments. Further specifics on the computation of the true value functions using Monte Carlo and details on network architectures are provided in Section B.2.
@@
In our study, we also integrated Prioritized Experience Replay (PER) (Schaul et al., 2015) to evaluate the effectiveness of our algorithm. Unlike the standard Experience Replay Buffer which uniformly samples experiences, PER usually assigns priorities to samples based on the magnitude of the TD error in the Q-network. PER influences the weighting of samples during gradient updates without modifying the data collection process. In contrast, GVFExplorer adjusts the sampling strategy based on the total variance, defined as the expected cumulative discounted squared TD errors over time (refer Equation 6). GVFExplorer proactively uses future errors to adapt the sampling strategy. Whereas, PER focuses on one-step errors and only alters gradient updates without affecting the sampling strategy. Our method differs from PER in its independence from the replay buffer size ( or example, online learning in tabular experiments), offering flexibility with limited memory, while PER is known for diminishing performance with smaller buffer sizes.
However, with further analysis, we found that PER is compatible with GVFExplorer, and in fact enhances its effectiveness. We use the absolute sum of TD errors across multiple GVF $Q$ functions as a priority metric for PER in all the baselines, including GVFExplorer. We also conducted experimental trials using TD error of variance function to set the priority in GVFExplorer, but the results are not as good as the one obtained with priority on $Q$ function’s TD error. In Figure 4, we present the MSE for both standard experience replay (solid lines) and PER (dotted lines) for all the algorithms. Interestingly PER reduces the MSE for most of the methods, but its integration with GVFExplorer offers a notable improvement. For the SR baseline, where the TD error in SR predictions is used as priority due to lack of a separate Q-network, we noted a substantial performance degradation. This suggests that the non-stationarity in the SRs’ TD errors under a dynamic behavior policy might lead PER to prioritize states that became less relevant under the current policy. The original SR work by McLeod et al. (2021) does not use PER in the experiments.
In Figure 5, we compare the absolute value errors for the baseline RoundRobin and GVFExplorer by discretizing the state space. Due to the objective of minimizing total MSE, we see smaller error magnitudes near the goals in GVFExplorer, indicating a focus on high-variance areas. Further insights into the variance estimation by GVFExplorer is shown in Section B.2. In Figure 6, GVFExplorer algorithm initially exhibits goal-directed behavior due to higher variability, resulting into faster cumulant propagation along other states. As the variance near the goals diminishes, the agent shifts its focus to other regions in descending order of variance, presenting a strategic data sampling approach. Finally, Table 1 summarizes the results highlighting the performance of various algorithms at $10^{6}$ steps.
@@
(b) MSE Left Goal
(c) MSE Right Goal
Figure 4: Averaged MSE in Continuous Env.: Compare the MSE metrics in baselines - RoundRobin, MixturePolicy, SR and GVFExplorer (averaged over $50$ runs with standard errors) for both standard Experience Replay Buffer (solid lines) and with Priority Experience Replay (PER) (dotted lines). GVFExplorer demonstrates lower MSE with both types of replay buffers. PER generally reduces MSE across all algorithms, except for SR.
(a) Avg $\hat{V}^{c_{i}}_{\pi_{i}}$ error(RoundRobin)
(b) $\hat{V}^{c_{1}}_{\pi_{1}}$ error(RoundRobin)
(c) $\hat{V}^{c_{2}}_{\pi_{2}}$ error(RoundRobin)
(d) Avg $\hat{V}^{c_{i}}_{\pi_{i}}$ error(GVFExplorer)
(e) $\hat{V}^{c_{1}}_{\pi_{1}}$ error(GVFExplorer)
(f) $\hat{V}^{c_{2}}_{\pi_{2}}$ error(GVFExplorer)
Table 1: Avg. MSE Summary for Cont. Env.: Averaged MSE across two GVFs for different algorithms in the continuous environment. GVFExplorer performance measured against others using standard and prioritized experience replay after $1\times 10^{6}$ learning steps. Note: Smaller MSE indicates better performance.
Standard Replay Buffer 21.7 18.25 16.78 5.19 69%
Prioritized Exp. Replay 112 14.7 11.62 3.87 66%
(a) RoundRobin $\tau_{1}$
(b) RoundRobin $\tau_{2}$
(c) GVFExplorer $\tau_{1}$
(d) GVFExplorer $\tau_{2}$
Figure 6: Sampled trajectories in Continuous Env: GVFExplorer method produces high variance-directed trajectories, depicting a smarter data-gathering strategy, as compared to RoundRobin which collects data according to target policies. Green dots show GVF goals and red depicts the start state.
@@ 9 Discussion
In this work, we addressed the problem of parallel evaluations of multiple GVFs, each conditioned on a given target policy and cumulant. We developed a method to adaptively learn a behavior policy that uses a single experience stream to estimate all GVF values in parallel. The behavior policy selects actions in proportion to the total variance of the return across the multiple GVFs. This guides the policy to explore less known areas, thereby accelerating information flow and minimizing environmental interactions. We theoretically proved that each behavior policy update step reduces the total prediction error. Our empirical analysis showed the scalability of GVFExplorer for a large number of GVFs in a tabular setting and its ability to adapt to non-linear function approximation.
@@
One notable drawback of GVFExplorer is the increased time complexity, due to simultaneously learning two networks - one for value and the other for variance estimation. Additionally, GVFExplorer has not been evaluated in environments with significant range variations in cumulant functions. Disparities in cumulant ranges could lead to very different variances and potentially result in oversampling areas with higher cumulant values. Some calibration across cumulants may be necessary in such cases.
In this work we focused on minimizing the total MSE, but other loss functions, such as weighted MSE could also be considered. However, weighted MSE requires a priori knowledge about the weighting of errors in different GVFs, which is not readily available. A potential future direction could be to use variance scales to automatically adjust these weights, in order to provide uniform MSE reduction across all GVFs. Looking ahead, we are interested in testing our approach with multi-dimensional cumulants and general state-dependent discount factors, as well as, extending the applicability of GVFExplorer to control settings where the target policies are unknown.
@@ Appendix A Proofs
We formulate Equation 3 as a Lagrangian equation below to solve for the optimal behavior policy $\mu^{*}$.
@@
We will compute the gradient of $M^{\mu}_{\pi_{i}}(s)$ in Equation 5 w.r.t to given $\mu(s,a)$. Here, $\rho(s,a)=\frac{\pi(a|s)}{\mu(a|s)}$ is IS weight. We expand $M^{\mu}_{\pi_{i}}(s)$ relation with $M^{\mu}_{\pi_{i}}(s,a)$ to derive the gradient,
@@
The final term $\nabla_{\mu}M^{\mu}_{\pi_{i}}(s,a)$ is difficult to estimate in the off-policy setting. Hence, we are omitting (IV part) from the above gradient, which is similar to Jain et al. (2021), where the gradient of $M(s,a)$ was omitted.
Solving for the Lagrangian Equation 8 further by substituting the (I part), and taking derivation of II & III part and using the (1) KKT condition.
@@
From (2)KKT condition, we know that either $\lambda_{s,a}=0$ or $\mu(a|s)=0$. Following the arguments of IS, support for $\mu(a|s)$ can only be 00 when the support for target policy $\pi(a|s)=0$. Solving for the case when support for target policy in non-zero, then let $\lambda_{s,a}=0$. We can simplify the gradient of Lagrangian in Equation 9,
@@
We know that the numerator is always positive (variance $M$ is positive), therefore $w_{s}<0$. Let $y_{s}=-w_{s}/d(s)$. From condition (5), we know that $\sum_{a}\mu(a|s)=1$. Using Equation 10 and summing over all the actions we get,
@@
Theorem 5.1 suggests, for any given $\mu_{k}$ behavior policy, the next successive approximation $\mu_{k+1}$ minimizes the objective function Equation 3, i.e.,
@@
$=\min_{\mu}\sum_{i}\sum_{s}d(s)\underbrace{\sum_{a}\mu(a|s)\dfrac{\pi_{i}(a|s)^{2}}{\mu(a|s)^{2}}M^{\mu_{k}}_{\pi_{i}}(s,a)}_{=\text{I}}.$
We will omit writing $d(s)$ and assume that $s\sim d(s)$. Further, we will use the notation $\rho^{i}_{k}(s,a)=\dfrac{\pi_{i}(a|s)}{\mu_{k}(a|s)}$ for ease of writing. From Equation 11, we can establish the relation,
@@
Now, we will use Equation 12 relation to further simplify the equation and establish that variance decreases with every update step $k$. We will use the notation $\rho_{t:t+n}=\Pi_{l=0}^{n}\rho_{t+l}$ to denote the products.
@@
Following Lemma A.3, the inverse $(I-\gamma^{2}\mathbf{\bar{P}_{\mu}})^{-1}$ exists if spectral radius $\textit{sr}(\gamma^{2}\mathbf{\bar{P}_{\mu}})<1$. Further, from Lemma A.2, we know that for any given matrix $\mathbf{A}$, spectral radius satisfies, $\textit{sr}(\mathbf{A})\leq\|A\|$. Hence, using the above two lemmas, we can express,
@@ Appendix B Experiments
We consider the two target policies with four cardinal directions left (L), right (R), up (U) and down (D) for the tabular and non-linear function approximation environments. These policies are specified as follows for every state $s\in\mathcal{S}$:
@@
B.1 Tabular Environment
We consider a tabular $20\times 20$ grid environment with stochastic dynamics. We use the target policies in Section B.1 for the different experimental settings. The Table 2 summarizes the averaged MSE with the same $2\times 10^{6}$ samples for different experimental settings.
@@ B.1 Tabular Environment
Hyperparameter Tuning:
In our experiments, we use linearly decaying learning rates that starts with initial value of $1.0$ and gradually decreased to an optimized minimum value within $500K$ steps of environmental interactions. We used different learning rates for value and variance function in GVFExplorer. The minimum learning rate parameter was swept within $\{0.1,0.25,0.5,0.8,0.9,0.95\}$ for both value and variance function. The optimal minimum learning rate was determined based on the one achieving the lowest average Mean Squared Error (MSE) after $800K$ sample interactions. This hyperparameter tuning approach was consistently applied for all algorithms including baselines. Figure 7 shows the sensitivity analysis of varying learning rates for value functions (all baselines) and variance functions (our method) with the averaged MSE performance in experiment Setting 1 with distinct target policies and same cumulant. The learning rate resulting in the lowest MSE was selected as optimal. In Table 3, we show the optimal hyperparameters for the four experimental settings discussed in the paper earlier (refer Experimental Settings in Section 8).
Baselines.
(2) MixturePolicy, UniformPolicy are also evaluated using Expected Sarsa. (3)SR: (McLeod et al., 2021) Used a summation of weight change in SR and reward weights to get the intrinsic reward for behavior policy. Further Expected Sarsa is used over this intrinsic reward to learn the Q value function for behavior policy. Following the original work of McLeod et al. (2021), we use greedy policy over this behavior’s $Q$ function to update the behavior policy. We use same learning rates for SR, reward weights and behavior policy Q function learning. (4)BPS: (Hanna et al., 2017) Use a Reinforce style estimator to learn $IS(\tau,\pi)=G(\tau)\Pi_{t=1}^{T}\dfrac{\pi(a_{t}|s_{t})}{\mu(a_{t}|s_{t})}$, as mentioned in the original paper. Since, the original work is only about single policy evaluation, we extended for multiple GVFs by updating behavior policy as summation over $\sum_{i}IS(\tau,\pi_{i})$. The behavior policy weight $\theta$ is updated as:
(a) RoundRobin
(b) SR
(c) BPS
(d) GVFExplorer
Figure 7: Impact of Learning Rate on Averaged MSE in Setting 1: Demonstrate the effect of changing minimum value of learning rate on the averaged MSE (performance averaged over $10$ runs) across GVFs in Setting 1. The optimal hyperparameter is selected based on the least MSE in these plots. LR_Q: value learning rate, LR_M: variance learning rate.
Experiment with Semi-greedy $\pi$ in 20x20 Grid :
In Setting 2, we evaluated semi-greedy target policies with distinct cumulants, $(\pi_{1},c_{1})$ and $(\pi_{2},c_{2})$ within a 20x20 grid. The target policies are designed with a bias towards top-left and top-right goals respectively,
$\pi_{1}(s)$ =$=\{L:0.4,R:0.1,U:0.4,D:0.1\}\forall s\in\mathcal{S}$ (15)
$\pi_{2}(s)$ =$=\{L:0.1,R:0.4,U:0.4,D:0.1\}\forall s\in\mathcal{S}.$
We keep the same cumulants as in previous experiments, $c_{1}=\mathcal{N}(\mu=100,\sigma=5)$ on top-left corner and $c_{2}=\mathcal{N}(\mu=50,\sigma=5)$ on top-right corner only, with 00 cumulant everywhere else. In Figure 8, we compare the average MSE performance, where GVFExplorer exhibits comparable MSE to RoundRobin baseline but requires more samples to converge. This outcome can be attributed to the near-greedy nature of the target policies, which inherently guides RoundRobin along goal directed trajectories, enabling it to achieve nearly accurate predictions with fewer samples. The optimal hyperparameters for RoundRobin, UniformPolicy and MixturePolicy is $\alpha_{Q}=0.95$. We used $\alpha_{q}=0.5,\alpha_{M}=0.8$ for ours GVFExplorer. Another baseline SR has $\alpha_{Q}=0.8$ and BPS as $\alpha_{Q}=0.9$ as optimal hyperparameters.
(a) Averaged MSE
(b) $\text{MSE}_{1},\text{MSE}_{2}$
Figure 8: 2 GVFs with distinct greedy target policy-cumulant pairs in 20x20 Grid: Analysis of MSE in value prediction for 2 semi-greedy target policies with different cumulant $(\pi_{1},c_{1}),(\pi_{2},c_{2})$ averaged over $25$ runs. We show results for baselines – RoundRobin, MixturePolicy, UniformPolicy, SR, BPS and GVFExplorer. (a) Averaged MSE, that is, $\mathbb{E}_{i}[\left(V_{\pi_{i}}^{c}(s)-\hat{V}_{\pi_{i}}^{c}(s)\right)^{2}]$. (b) $\text{MSE}_{1},\text{MSE}_{2}$. We observe a slower convergence of GVFExplorer as compared to baselines like RoundRobin, MixturePolicy due to target policies being semi-greedy. Therefore, baselines lead to goal-directed policies, hence converging faster.
Experiment with high number of GVF predictions, Setting 3:
Figure 9: $10$ different cumulants in high number of GVF prediction experiment in 20x20 grid. The color depict the cumulant empirical value.
In this setting, we considered $4$ target policies in the four cardinal directions, namely:
$\pi_{N}(s)$ =$=\{L:0.1,R:0.1,U:0.7,D:0.1\}\forall s\in\mathcal{S}$
$\pi_{E}(s)$ =$=\{L:0.1,R:0.7,U:0.1,D:0.1\}\forall s\in\mathcal{S}$
$\pi_{S}(s)$ =$=\{L:0.1,R:0.1,U:0.1,D:0.7\}\forall s\in\mathcal{S}$
$\pi_{W}(s)$ =$=\{L:0.7,R:0.1,U:0.1,D:0.1\}\forall s\in\mathcal{S}.$
We considered $10$ different cumulants, each with a single goal of constant reward at states uniformly choose from $\mathcal{S}$. The cumulant value is uniformly choosen between $\{50,100\}$ and is depicted in Figure 9. In total, the task is to predict $40$ GVF values. The experimental results are described in the main paper.
@@
Exp. Settings Setting 1 Setting 2 Setting 3 Setting 4

### v2

@@ 7.2 Continuous State Environment with Non-Linear Function Approximation
We use a continuous grid environment that extends the tabular experiments to a continuous state space (similar to McLeod et al. (2021)) and four discrete actions. For Non-Linear Function Approximation, we consider two distinct GVFs with distractor cumulants. An Experience Replay Buffer with a capacity of 25K and a batch size of 64 is used for all experiments. Further details on computing true value functions using Monte Carlo and network architectures are in Sec. B.4.
@@
We integrate PER (Schaul et al., 2015) to evaluate the effectiveness of our algorithm. Unlike the standard Experience Replay Buffer, which uniformly samples experiences, PER assigns priorities based on the TD error magnitude in the Q-network. PER and GVFExplorer are complementary approaches: PER re-weights the collected data in replay buffer based on the priority, while GVFExplorer adjusts the behavior policy to influence data collection.
Combining PER with GVFExplorer drastically lowers MSE compared to other baselines (even when compared to all baselines + PER). We use the absolute sum of TD errors across multiple GVF Q-functions as a priority metric for PER in all baselines, including GVFExplorer. Placing the priority on the TD error of the variance function in GVFExplorer yields less favorable results compared to priority on Q-function’s TD error. In Fig. 3, we present the MSE for both standard experience replay (solid lines) and PER (dotted lines) for all algorithms. PER generally reduces MSE, but its integration with GVFExplorer shows much lower MSE. This is likely as GVFExplorer could over-sample high variance return samples, causing a skewed buffer distribution. PER’s non-uniform sampling maintains a balanced data distribution, which helps in stringent MSE reduction. For the SR baseline, using the TD error in SR predictions as a priority for PER led to performance degradation, suggesting non-stationarity in SRs’ TD errors might mislead PER to prioritize less relevant states under the current policy. The original SR work by McLeod et al. (2021) does not use PER in the experiments. For PER scenario, we qualitatively compare the absolute value error for baseline RoundRobin and GVFExplorer by discretizing the state space in Figs. 3(b) and 3(c) and observe that our algorithms results in smaller value prediction error. Further insights into the variance estimation by GVFExplorer is shown in Figs. 16 and 15 (Sec. B.4). Table 3((Sec. B.4) summarizes the results highlighting the performance of various algorithms.
@@
(b) Value error (RoundRobin)
(c) Value error (GVFExplorer)
Figure 3: Non-Linear Function Approximation: (a) Averaged MSE over $50$ runs with standard error using Experience Replay Buffer (solid lines) and PER (dotted lines). GVFExplorer show lower MSE with both buffers. PER generally reduces MSE across all algorithms except SR. Log-scale absolute value error for RoundRobin (b) and GVFExplorer (c); GVFExplorer achieves smaller errors (vibrant colors represent higher values).
7.3 Mujoco Environments with Continuous State-Actions Tasks
We use DM-Control (Tassa et al., 2018) based continuous state-action tasks to experiment with Mujoco environments, Walker and Cheetah domain. To expand the proposed method to continuous action environments, any policy-gradient (PG) based algorithm can be used. In our experiments, we use Soft Actor-Critic (SAC) algorithm (Haarnoja et al., 2018) as a base PG method to incorporate the proposed variance-minimization objective.
A separate network for variance estimation is added to SAC. Further implementation details are provided in Sec. B.5. To experiment in Walker environment, we use two GVF tasks, namely walk and flip. Similarly, for Cheetah environment, we use walk and run GVF tasks. We also added KL regularize between the learned behavior policy and the given GVFs target policies to prevent divergence. We use MC to compute the true Q-value GVF estimates and compare the MSE between these MC values and the Q-critic network. We use the same Q-critic architecture for the baseline algorithms – UniformPolicy and RoundRobin – for fair comparison. In Fig. 4 we observe that GVFExplorer reduces MSE faster than the baselines.
(a) Averaged MSE in Walker
(b) Averaged MSE in Cheetah
Figure 4: MSE in Mujoco: Averaged MSE over $5$ runs with standard error in Mujoco environment with continuous state-actions for (a)Walker and (b)Cheetah domains for GVFExplorer, UniformPolicy and RoundRobin. GVFExplorer consistently lowers averaged MSE as compared to the baselines.
@@ 8 Discussion
We addressed the problem of parallel evaluations of multiple GVFs, each conditioned on a given target policy and cumulant. We developed a method to adaptively learn a behavior policy that uses a single experience stream to estimate all GVF values in parallel. The resulting behavior policy update selects the actions in proportion to the total variance of the return across GVFs. This guides the policy to frequently explore less understood areas (high variance in return), which helps to better estimate the mean return with fewer samples. Therefore, our approach lowers the overall MSE in GVF predictions and uses fewer interactions. We theoretically proved that each behavior policy update reduces the total prediction error. Empirically, we showed that GVFExplorer scales effectively with an increasing number of distinct GVFs, robustly handles non-stationary cumulants in a tabular setting, and adapts well to non-linear function approximation. Further, we also demonstrated the performance in complex continuous state-action Mujoco environments by showing that GVFExplorer can be combined effectively with any existing policy-gradient methods.
@@
One notable drawback of GVFExplorer is the increased time complexity, due to simultaneously learning two networks for value and variance estimation respectively. Additionally, GVFExplorer has not been evaluated in environments with significant difference in the cumulant value range. Such disparities could lead to varying variances, potentially resulting in oversampling areas with higher cumulant values. Calibration across cumulants may be necessary in these cases.
In this work, we focused on minimizing the total MSE, but other loss functions, such as weighted MSE could also be considered. However, weighted MSE requires prior knowledge about the weighting of errors in different GVFs, which is not readily available. A potential future direction could be to use variance scales to automatically adjust these weights to provide uniform MSE reduction across all GVFs. Looking ahead, we are interested in testing our approach with multi-dimensional cumulants and general state-dependent discount factors, as well as, extending the applicability of GVFExplorer to control settings where the target policies are unknown.
Acknowledgments and Disclosure of Funding
We are grateful to the anonymous reviewers for their valuable feedback. We also extend our thanks to Nishanth Anand, Kshitij Jain, Pierre-Luc Bacon, and Subhojyoti Mukherjee for their insightful suggestions.
@@ A.1 Behavior Policy Update Theorems
We formulate Eq. 2 as a Lagrangian equation below to solve for the optimal behavior policy $\mu^{*}$.
@@
We will compute the gradient of $M^{\mu}_{\pi_{i}}(s)$ in Eq. 4 w.r.t to given $\mu(s,a)$. Here, $\rho(s,a)=\frac{\pi(a|s)}{\mu(a|s)}$ is IS weight. We expand $M^{\mu}_{\pi_{i}}(s)$ relation with $M^{\mu}_{\pi_{i}}(s,a)$ to derive the gradient,
@@
The final term $\nabla_{\mu}M^{\mu}_{\pi_{i}}(s,a)$ is difficult to estimate in an iterative off-policy setting. Hence, we are omitting (IV part) from the above gradient, which is similar to Degris et al. (2012)[Sec 2.2], where the gradient of $Q(s,a)$ was omitted while deriving the policy update.
Solving for the Lagrangian Eq. 7 further by substituting the (I part), and taking derivation of II & III part and using the (1) KKT condition.
@@
From (2)KKT condition, we know that either $\lambda_{s,a}=0$ or $\mu(a|s)=0$. Following the arguments of IS, support for $\mu(a|s)$ can only be 00 when the support for target policy $\pi(a|s)=0$. Solving for the case when support for target policy in non-zero, then let $\lambda_{s,a}=0$. We can simplify the gradient of Lagrangian in Eq. 8,
@@
We know that the numerator is always positive (variance $M$ is positive), therefore $w_{s}<0$. Let $y_{s}=-w_{s}/d(s)$. From condition (5), we know that $\sum_{a}\mu(a|s)=1$. Using Eq. 9 and summing over all the actions we get,
@@
Theorem 4.1 suggests, for any given $\mu_{k}$ behavior policy, the next successive approximation $\mu_{k+1}$ minimizes the objective function Eq. 2, i.e.,
@@
$=\min_{\mu}\sum_{i}\sum_{s}d(s)\underbrace{\sum_{a}\mu(a|s)\dfrac{\pi_{i}(a|s)^{2}}{\mu(a|s)^{2}}M^{\mu_{k}}_{\pi_{i}}(s,a)}_{=M^{\mu_{k}}_{\pi_{i}}(s)}.$
We will omit writing $d(s)$ and assume that $s\sim d(s)$. Further, we will use the notation $\rho^{i}_{k}(s,a)=\dfrac{\pi_{i}(a|s)}{\mu_{k}(a|s)}$ for ease of writing. From Eq. 10, we can establish the relation,
@@
Now, we will use Eq. 11 relation to further simplify the equation and establish that variance decreases with every update step $k$. We will use the notation $\rho_{t:t+n}=\Pi_{l=0}^{n}\rho_{t+l}$ to denote the products.
@@ A.2 When does Variance Function Exists?
Following Lemma A.3, the existence of $M$ hinges on the existence of inverse $(I-\gamma^{2}\mathbf{\bar{P}_{\mu}})^{-1}$. Further, $(I-\gamma^{2}\mathbf{\bar{P}_{\mu}})^{-1}$ exists if spectral radius $\textit{sr}(\gamma^{2}\mathbf{\bar{P}_{\mu}})<1$. Further, from Lemma A.2, we know that for any given matrix $\mathbf{A}$, spectral radius satisfies, $\textit{sr}(\mathbf{A})\leq\|A\|$. Hence, using the above two lemmas, we can express,
@@ Appendix B Experiments
This section provides detail about the experiments in the main paper as well as additional experiments in the Appendix. All the experiments require less than $1GB$ of memory and have used combined compute less than total 4 CPU months and 1 GPU month.
For all the experiments, we consider the two target policies with four cardinal directions left (L), right (R), up (U) and down (D) for the tabular and non-linear function approximation environments. These policies are specified as follows for every state $s\in\mathcal{S}$:
@@
B.1 Baselines
(2) MixturePolicy, UniformPolicy are also evaluated using Expected Sarsa. MixturePolicy is defined as,
$\mu_{\text{Mixture}}(a|s)=\dfrac{\sum_{i=1}^{\mathcal{N}}\pi_{i}(a|s)}{\sum_{a^{\prime}}\sum_{i=1}^{\mathcal{N}}\pi_{i}(a^{\prime}|s)}.$
(3) SR: Based on (McLeod et al., 2021), SR uses the summation of weight changes in the Successor Representation (SR) and reward weights to obtain the intrinsic reward for behavior policy updates. We use Expected Sarsa to learn both the SR and the Q-value function from the intrinsic reward. The behavior policy is generated using a Boltzmann policy over the learned Q function, as it empirically performs better than a greedy policy. We apply simple TD Expected Sarsa updates instead of Emphatic TD($\lambda$) as shown in Algo 2 in McLeod et al. (2021). The learning rates for SR, reward weights, and behavior policy Q function are kept the same.
(4)BPS: (Hanna et al., 2017) Use a Reinforce style estimator to learn $IS(\tau,\pi)=G(\tau)\Pi_{t=1}^{T}\frac{\pi(a_{t}|s_{t})}{\mu(a_{t}|s_{t})}$, as mentioned in the original paper. Since, the original work is only about single policy evaluation, we extended for multiple GVFs by updating behavior policy as summation over $\sum_{i}IS(\tau,\pi_{i})$. The behavior policy weight $\theta$ is updated as:
B.2 Types of Cumulants
Figure 5: Visual representation of cumulants.
We consider three different types of cumulants similar to McLeod et al. (2021) as shown in Fig. 5:
  • Constant: Fixed value cumulant, $c_{i}^{t}=c_{i}$
  • Distractor: Stationary cumulant with reward drawn from normal distribution with fixed mean and variance, $c_{i}^{t}=\mathcal{N}(\mu_{i},\sigma_{i})$
  • Drifter: Non-Stationary cumulant whose value change over time, $c_{i}^{t}=c_{i}^{t-1}+\mathcal{N}(\mu_{i},\sigma_{i}),c_{i}^{0}=100.$
B.3 Tabular Experiments
We consider a tabular $20\times 20$ grid environment with stochastic dynamics. We use the above target policies for the different experimental settings. The Table 1 summarizes the averaged MSE with the same $2\times 10^{6}$ samples for different experimental settings.
@@ B.3 Tabular Experiments
Exp. Settings distinct policies identical cumulants distinct policies distinct cumulants large scale $40$ GVF eval non-linear func approx non-stationary cumulant in FR